-x+4y=20
how to solve this equation in slope intercept form.
then find the slope y-intercept

Answer by sudhanshu_kmr(1152)


-x+4y=20
=> 4y = x + 20
=> y = x/4 + 20/4
=> y = x/4 + 5 (slope intercept form)

slope = 1/4, y-intercept = 5